OpenStudy (rootbeer003):

Which point is on the graph of y=3/X? (6, 2) (6, ) (0, 3) (3, 0)

OpenStudy (rootbeer003):

rIGHT BY 6 IS 1/2

jimthompson5910 (jim_thompson5910):

Let's say we replace x with 3 if we do that, we get... y = 3/x y = 3/3 ... replaced x with 3 y = 1 so when x = 3, the value of y is y = 1 Together they form the ordered pair (3,1) So the point (3,1) is on the curve but (3,0) is NOT on the curve

jimthompson5910 (jim_thompson5910):

Now let's try x = 1 y = 3/x y = 3/1 ... replace every x with 1 y = 3 So the point (x,y) = (1,3) is also on the curve

OpenStudy (rootbeer003):

hmm so its ( 3,0)?

jimthompson5910 (jim_thompson5910):

I just showed how (3,1) was on the graph. We cannot have (3,0) on the graph because the input x = 3 leads to the output y = 1 any given input leads to exactly one output

jimthompson5910 (jim_thompson5910):

make sense?